# A Survey of Pathology Foundation Model: Progress and Future Directions

### v1

@@ 3.3 Model Design
By model design, we refer to the following three aspects that are vital to model performance: architecture, number of parameters (# params.), scale. The scale of a model is directly determined by its number of parameters. Through quantization of the number of parameter, we establish a hierarchical scale system, facilitating standardized cross-architectural comparisons and enabling informed model selection for practical implementations. The taxonomy of PFMs along this dimension is presented in the Model Design column of Table 1.
Architecture is a pivotal determinant of PFM capabilities. Architectures adopted by PFMs can be categorized based on scope: extractor architecture and aggregator architecture. The extractor architecture encompasses two categories: 1) CNN-based architecture, ResNet, and 2) transformer-based architecture, including ViT Dosovitskiy et al. (2021), CNN-integrated Swin Transformer Wang et al. (2022), BEiTv2 Peng et al. (2022), FlexiViT Beyer et al. (2023), and BEiT-3 Wang et al. (2023), a multimodal one. The aggregator architecture comprises two categories: the ABMIL family Ilse et al. (2018); Ding et al. (2024) and Perceiver Jaegle et al. (2021). For an architectural overview of each method, we refer readers to the Architecture column in Table 1.
Scale can be derived through the number of parameters. To facilitate cross-architectural comparisons, we establish a quantization framework based on the ViT architecture, which serves as the predominant backbone across our surveyed literature. The classification includes seven categories: extra small (XS, 2.78M), Small (S, 21.7M), Base (B, 86.3M), Large (L, 307M), Huge (H, 632M), giant (g, 1.13B), and Giant (G, 1.9B). The notation ViT-B/16 indicates a ViT Base model with patch size 16. For statistics for each method, we refer readers to the # Params. and Scale columns in Table 1.
Our analysis reveals several patterns in model architectures and scaling: 1) ABMIL-derived methods demonstrate clear dominance in aggregator architectures, while the ViT family predominates in extractor architectures, both are transformer-based architectures; 2) The majority of methods utilize ViT-L as their primary backbone. Due to computational resource constraints, researchers often develop complementary smaller-scale variants (ViT-S or ViT-B) alongside their primary models, while some approaches specifically target efficiency through smaller architectures; 3) ViT-L is a popular scale for extractors, whereas ViT-XS is the primary choice for aggregators, even other than ViT-S. This substantial disparity in parameter counts between extractors and aggregators, despite their similar training data scale, suggests a potential data-model scale mismatch that warrants further investigation; 4) a clear trend toward larger model scales is observed: while earlier approaches frequently employed ViT-B, recent methods have increasingly standardized on ViT-L, with some extending to even larger variants such as ViT-H/g/G.
@@ 4 Evaluation Tasks for the Foundation Model
Development and evaluation constitute the two fundamental pillars of PFM. The evaluation tasks of PFMs can be systematically categorized into four aspects: 1) slide-level tasks; 2) patch-level tasks; 3) multimodal tasks; and 4) biological tasks. A comparative analysis of these evaluation tasks is presented in Table 3, providing practitioners with comprehensive criteria for model selection based on real-world applications.
@@
Multimodal Tasks are tasks that evaluate multimodal capabilities of PFMs. These tasks encompass cross-modal retrieval, i.e., image-to-text (I2T) and text-to-image (T2I) retrieval, report generation (RG), and visual question answering (VQA). RG in our survey includes both RG and image captioning, distinguished by their input: utilizes WSIs to generate clinical documentation, while captioning produces concise descriptions from patches. The increasing emphasis on these tasks reflects the clinical reality that pathologists integrate multimodal data in decision-making process.
Biological Tasks focus on biomarker detection, including genetic alteration (GA) and molecular prediction (MP). Genetic alteration includes both mutation prediction and genetic alteration, as both predicts gene mutation status. Molecular prediction targets the prediction of molecular subtypes at the gene expression level, representing a distinct biomarker from genetic alteration. While these tasks can be fundamentally categorized as classification problems at either slide or patch level, their clinical applications and biological implications warrant their classification as a separate analytical category. One recently-proposed task is molecular prompting Vaidya et al. (2025), which aims to perform clinical tasks with canonical molecular profiles without requiring any task-specific model development in a similar manner to text prompting.
@@ 5.1 Foundation Model Development
End-to-end Pretraining is critical to achieve optimal performance for PFMs. Current PFMs adopt a two-stage pretraining paradigm: extractors are trained independently, followed by aggregator with the extractor frozen. Evidence suggests this complicates optimization, highlighting the need for end-to-end pretraining of PFMs, which poses significant challenges in CPath, as transitioning away from MIL requires developing extremely sophisticated and efficient architectures and algorithms capable of simultaneously integrating local and global pathology information for gigapixel images.
@@
Data-Model Scalability is a critical direction, as performance improvements continue to demonstrate logarithmic and sub-logarithmic scaling with model and data volume, respectively, without yet reaching a clear plateau. This domain presents four sub-directions: 1) examining the relative importance of WSI and patch quantity, particularly when considering diversity, a complex concept yet widely acknowledged as an indicator of high-quality data; 2) exploring efficient algorithms, due to the rapid expansion in both datasets and models, evident in transitions from CONCH (ViT-B) to CONCHv1.5 (ViT-L) and from UNI (ViT-L) to UNI2 (ViT-H); 3) addressing data-model scale mismatch problem for aggregator, detailed in Sec. 3.1; and 4) optimizing model scale, since the giant model size poses substantial deployment challenges in both hospital environments and academic settings.
@@
Federated Learning with Efficiency is essential for addressing the challenges associated with collecting massive-scale datasets across multiple institutions while preserving patient privacy, as few institutions can feasibly collect WSIs at million scale alone. However, current research in this area remains limited; for instance, HistoFL has demonstrated improved performance, yet this benefit come at the cost of significantly increased computational overhead. As PFMs continue to grow in size, scaling federated learning further exacerbates these challenges. Consequently, there is an urgent need to develop more efficient, privacy-protected frameworks in such large-scale cross-institution collaboration settings.
@@
Model Robustness addresses critical challenges in multi-institutional data curation. The acquisition of data from diverse institutions inevitably introduces technical heterogeneity across scanning equipment specifications, image magnification levels, and staining protocols, resulting in significant data variations. These technical disparities can induce training instability, even worse with federated learning. Consequently, the development of more robust pretraining algorithms for PFMs is a critical research imperative.
@@
RAG-enhanced Pathology VLM is an trending paradigm worth investigation. Contemporary trends in Large Language Models (LLMs), coupled with the prevalence of BERT-based architectures in current multimodal PFMs, suggest the potential utility of integrating advanced LLMs with ViT architectures. Furthermore, given the demonstrated efficacy of Retrieval-Augmented Generation (RAG) in LLMs and the critical need for domain-specific expertise in pathology, RAG methodology offers promising avenues for enhancing representation learning in pathology VLMs. This approach transcends the limitations of existing methods such as RudolfV, which relies primarily on clustering techniques for pathologist knowledge integration, providing a potentially more sophisticated framework for incorporating domain expertise.
@@ 5.2 Foundation Model Utilization
Model Maintenance constitutes a critical research domain in the context of PFMs, given the substantial computational resources required for their initial training. The potential diminishment of model performance due to novel diseases, tissue heterogeneity, or technological advancements necessitates efficient maintenance strategies to preserve model utility. Continual learning represent a promising approach for maintaining PFM effectiveness, as they circumvent the necessity for model retraining by learning on newly observed instances. This approach significantly reduces computational overhead while ensuring the model remains current with evolving clinical and technological developments.
@@ 6 Conclusion
This survey presents a systematic analysis of Pathology Foundation Models through our proposed taxonomy and comprehensive evaluation framework. While PFMs demonstrate significant advances, critical technical challenges merits further investigation. We delineate key directions that are worth exploring and might be instrumental in advancing both the theoretical foundations and practical applications of PFMs.

### v2

@@ 3.3 Model Design
The model design refers to the following three aspects that are vital to model performance: architecture, number of parameters (# params.), scale. The scale of a model is directly determined by its number of parameters. Through quantization of the number of parameters, we establish a hierarchical scale system, facilitating standardized cross-architectural comparisons and enabling informed model selection for practical implementations. The taxonomy of PFMs along this dimension is presented in the Model Design column of Table 1.
Architecture is a pivotal determinant of PFM capabilities. Architectures adopted by PFMs can be categorized based on scope: extractor architecture and aggregator architecture. The extractor architecture encompasses two categories: 1) CNN-based architecture such as ResNet, and 2) transformer-based architecture, including ViT Dosovitskiy et al. (2021), CNN-integrated Swin Transformer Wang et al. (2022), BEiTv2 Peng et al. (2022), FlexiViT Beyer et al. (2023), and a multimodal BEiT-3 Wang et al. (2023). The aggregator architecture comprises two categories: ABMIL family Ilse et al. (2018); Ding et al. (2024) and Perceiver Jaegle et al. (2021). For an architectural overview of each method, we refer readers to the Architecture column in Table 1.
Scale can be derived through the number of parameters. To facilitate cross-architectural comparisons, we establish a quantization framework based on the ViT architecture, which serves as the predominant backbone across our surveyed literature. The classification includes seven categories: extra small (XS, 2.78M), Small (S, 21.7M), Base (B, 86.3M), Large (L, 307M), Huge (H, 632M), giant (g, 1.13B), and Giant (G, 1.9B). The notation ViT-B/16 indicates a ViT Base model with patch size 16. For statistics of each method, we refer readers to the # Params. and Scale columns in Table 1.
Our analysis reveals several patterns in model architectures and scaling: 1) ABMIL-derived methods demonstrate clear dominance in aggregator architectures, while the ViT family predominates in extractor architectures, which are transformer-based architectures; 2) The majority of methods utilize ViT-L as their primary backbone. Due to computational resource constraints, researchers often develop complementary smaller-scale variants (ViT-S or ViT-B) alongside their primary models, while some approaches specifically target efficiency through smaller architectures; 3) ViT-L is a popular scale for extractors, whereas ViT-XS is the primary choice for aggregators, over ViT-S. This substantial disparity in parameter counts between extractors and aggregators, despite their similar training data scale, suggests a potential data-model scale mismatch that warrants further investigation; 4) a clear trend toward larger model scales is observed: while earlier approaches frequently employed ViT-B, recent methods have increasingly standardized on ViT-L, with some extending to even larger variants such as ViT-H/g/G.
@@ 4 Evaluation Tasks for the Foundation Model
Development and evaluation constitute the two fundamental pillars of PFMs. The evaluation tasks of PFMs can be systematically categorized into four aspects: 1) slide-level tasks; 2) patch-level tasks; 3) multimodal tasks; and 4) biological tasks. A comparative analysis of these evaluation tasks is presented in Table 3, providing practitioners with comprehensive criteria for model selection based on real-world applications.
@@
Multimodal Tasks are tasks that evaluate multimodal capabilities of PFMs. These tasks encompass cross-modal retrieval, i.e., image-to-text (I2T) and text-to-image (T2I) retrieval, report generation (RG), and visual question answering (VQA). RG in our survey includes both RG and image captioning, distinguished by their input: RG utilizes WSIs to generate clinical documentation, while captioning produces concise descriptions from patches. The increasing emphasis on these tasks reflects the clinical reality that pathologists integrate multimodal data in the decision-making process.
Biological Tasks focus on biomarker detection, including genetic alteration (GA) and molecular prediction (MP). Genetic alteration includes both mutation prediction and genetic alteration, as both predict gene mutation status. Molecular prediction targets the prediction of molecular subtypes at the gene expression level, representing a distinct biomarker from genetic alteration. While these tasks can be fundamentally categorized as classification problems at either slide or patch level, their clinical applications and biological implications warrant their classification as a separate analytical category. One recently-proposed task is molecular prompting Vaidya et al. (2025), which aims to perform clinical tasks with canonical molecular profiles without requiring any task-specific model development in a similar manner to text prompting.
@@ 5.1 Foundation Model Development
End-to-end Pretraining is critical to achieve optimal performance for PFMs. Current PFMs adopt a two-stage pretraining paradigm: extractors are trained independently, followed by the aggregator with the extractor frozen. Evidence suggests this complicates optimization, highlighting the need for end-to-end pretraining of PFMs, which poses significant challenges in CPath, as transitioning away from MIL requires developing extremely sophisticated and efficient architectures and algorithms capable of simultaneously integrating local and global pathology information for gigapixel images.
@@
Data-Model Scalability is a critical direction, as performance improvements continue to demonstrate logarithmic and sub-logarithmic scaling with model and data volume, respectively, without yet reaching a clear plateau. This domain presents four sub-directions: 1) examining the relative importance of WSI and patch quantity, particularly when considering diversity, a complex concept that is yet widely acknowledged as an indicator of high-quality data; 2) exploring efficient algorithms, due to the rapid expansion in both datasets and models, evident in transitions from CONCH (ViT-B) to CONCHv1.5 (ViT-L) and from UNI (ViT-L) to UNI2 (ViT-H); 3) addressing the data-model scale mismatch problem for the aggregator, detailed in Sec. 3.1; and 4) optimizing model scale, since the giant model size poses substantial deployment challenges in both hospital and academic settings.
@@
Federated Learning with Efficiency is essential for addressing the challenges associated with collecting massive-scale datasets across multiple institutions while preserving patient privacy, as few institutions can feasibly collect WSIs at the million-scale alone. However, current research in this area remains limited; for instance, HistoFL Lu et al. (2022) has demonstrated improved performance, yet this benefit comes at the cost of significantly increased computational overhead. As PFMs continue to grow in size, scaling federated learning further exacerbates these challenges. Consequently, there is an urgent need to develop more efficient, privacy-protected methods in such large-scale cross-institutional collaborations.
@@
Model Robustness addresses critical challenges in multi-institutional data curation. The acquisition of data from various sites inevitably introduces technical heterogeneity across scanning equipment specifications, image magnification levels, and staining protocols, resulting in significant data variations that embed site information de Jong et al. (2025). These disparities undermine training stability and model generalizability; recent work shows that most models encode site information more strongly than biological signals de Jong et al. (2025). These issues will be further exacerbated in federated learning under non-IID data distributions. Consequently, developing more robust algorithms and robustness evaluation metrics for PFMs is a critical research imperative.
@@
RAG-enhanced Pathology VLM is a trending paradigm worth investigating. Contemporary trends in Large Language Models (LLMs), such as Llama Grattafiori et al. (2024), with the prevalence of BERT-based architectures Devlin et al. (2019) in current multimodal PFMs suggest the potential utility of integrating LLMs with ViT architectures. Furthermore, given the demonstrated efficacy of Retrieval-Augmented Generation (RAG) Gao et al. (2023) in LLMs and the critical need for domain-specific expertise in pathology, RAG methodology offers promising directions for representation learning in pathology VLMs. This approach transcends the limitations of existing methods such as RudolfV, which relies primarily on clustering techniques for pathologist knowledge integration, providing a potentially more sophisticated framework for incorporating domain expertise.
@@ 5.2 Foundation Model Utilization
Model Maintenance constitutes a critical research domain in the context of PFMs, given the substantial computational resources required for their initial training. The potential diminishment of model performance due to novel diseases, tissue heterogeneity, or technological advancements necessitates efficient maintenance strategies to preserve model utility. Continual learning Wang et al. (2024a); Yu et al. (2024) represents a promising approach for maintaining PFM effectiveness, as it circumvents the necessity for model retraining by learning on newly observed instances. This approach significantly reduces the required computational overhead while ensuring the model remains current with the evolving clinical, disease, and technological developments.
@@ 6 Conclusion
This survey presents a systematic analysis of the current Pathology Foundation Models through our proposed hierarchical taxonomy and comprehensive evaluation framework. Although the PFMs demonstrate significant advances in computational pathology, critical technical challenges merit further investigation. We delineate key directions that are worth exploring and might be instrumental in advancing both the theoretical foundations and practical applications of PFMs.
Acknowledgments
The authors would like to sincerely thank Professor Irwin King from the Department of Computer Science and Engineering, the Chinese University of Hong Kong, for his active involvement in the conception and development of this paper. Due to the one-submission-per-author policy, his name could only appear in the acknowledgments. The work described in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 2410072, RGC R1015-23).